If I were born in 1978 and my child born in 2002. how old would I be when my child is half my age!!!
Answer by ankor@dixie-net.com(22740)

If I were born in 1978 and my child born in 2002.
how old would I be when my child is half my age!!!
:
This will be true when you are twice the difference in your ages
:
But step-by-step algebra method:
:
Let a = your age when your child is half your age
:
Given: 2002 - 1978 = 24 yrs difference in your ages
therefore
(a - 24) = your child's age
:
Simple equation to find a
a = 2(a-24)
a = 2a - 48
48 = 2a - a
48 = a
:
you will be 48 when you daughter is 24